# Chapter 0 Liquid-gas phase transition of nuclear matter

### v1

@@ 2 Liquid-gas transition and chemical freeze-out
High-energy collisions of nuclei trigger a sequence of events from compression, forming hot and possibly dense matter for a short time, through expansion and cooling to a stage at which finally non-interacting hadrons and nuclei emerge. This final stage is referred to as chemical freeze-out. It is explored by systematically tracing produced particles in such collisions and analysing the characteristic temperatures and chemical potentials at which these particles finally appear as decoupled from one another. In a $\mu-T$ phase diagram these chemical freeze-out points line up at temperatures far above and baryon chemical potentials well below the liquid-gas transition as presented in Figure 12.
The profile of empirical chemical freeze-out data in the phase diagram has sometimes tentatively been discussed as a possible extension towards finite $\mu$ of the crossover transition in QCD that has been established at high temperatures $(T\sim 150$ MeV) and vanishing baryon chemical potential. However, this freeze-out pattern and its position relative to the LGT critical line do not suggest such an interpretation. In fact the freeze-out data displayed in Fig. 12 turn out all to be located close to a curve of constant, very low baryon density (a small fraction of the nuclear equilibrium density $n_{0}$), that extends roughly from the LGT critical point upward in temperature and downward in chemical potential (Floerchinger et al., 2012). In the $\mu-T$ range covered by Fig. 12 there is no indication of a phase transition other than the LGT.
@@ 7 Summary
[Acknowledgment]  This work has been partially supported by the Excellence Cluster ORIGINS of the German Science Foundation (DFG).

### v2

@@ 2 Liquid-gas transition and chemical freeze-out
High-energy collisions of nuclei trigger a sequence of events from compression, forming hot and possibly dense matter for a short time, through expansion and cooling to a stage at which finally non-interacting hadrons and nuclei emerge. This final stage is referred to as chemical freeze-out. It is explored by systematically tracing produced particles in such collisions and analysing the characteristic temperatures and chemical potentials at which these particles finally appear as decoupled from one another. In a $\mu-T$ phase diagram these chemical freeze-out points line up at temperatures far above and baryon chemical potentials well below the liquid-gas transition (Braun-Munzinger et al., 2026, and refs. therein).
The systematics of empirical chemical freeze-out data in the phase diagram resembles the deconfinement and chiral crossover transition bands deduced and extrapolated from Lattice QCD. These bands extend over a range of low baryon chemical potentials and temperatures $160\,\text{MeV}\gtrsim T\gtrsim 140\,\text{MeV}$, starting from the crossover at $T\simeq 157$ MeV and zero chemical potential as it is widely agreed upon (Braun-Munzinger et al., 2026). Several chemical freeze-out points are also recorded at lower temperatures, $T<100$ MeV, and larger baryon chemical potential reaching out to $\mu\simeq 650-850$ MeV, as shown in Fig. 12. This data set, which one might envisage as continuing toward the LGT region further down in temperature, has sometimes been tentatively discussed as a possible extension of the QCD crossover transition to lower $T$ and larger $\mu$. However, this pattern and its position relative to the LGT critical line do not suggest such an interpretation. In fact the freeze-out data displayed in Fig. 12 turn out all to be located close to a curve of constant, very low baryon density (a small fraction of the nuclear equilibrium density $n_{0}$, see Floerchinger et al., 2012). It has thus been pointed out that no phase transition other than the LGT is expected in the range $T<100$ MeV and $0.65\,\text{GeV}\lesssim\mu\lesssim 0.95\,\text{GeV}$ covered by Figure 12.
@@ 7 Summary
[Acknowledgments]  This work has been partially supported by the Excellence Cluster ORIGINS of the German Science Foundation (DFG). Instructive comments by Peter Braun-Munzinger are gratefully acknowledged.